# Distill-SynthKG: Distilling Knowledge Graph Synthesis Workflow for Improved Coverage and Efficiency

### v1

@@ 5 Proposition-Entity Graph Retriever
We introduce a novel retriever based on Proposition-Entity Graph (Figure 2), designed for queries requiring multi-hop reasoning. Given a question, we first retrieve the top-M most relevant propositions from the knowledge graph using embedding similarity, narrowing the search space to a smaller subset of relevant information. In step 2, we construct a sub-graph consisting of these propositions and their linked entities, capturing the relations among the retrieved propositions. In step 3, we traverse the sub-graph starting from the entities mentioned in the question, and select only propositions within their N-hop neighborhood. This process filters out semantically similar but irrelevant propositions, ensuring that only those logically connected to the question entities are retained. We then include text chunks corresponding to the selected propositions within N-hop distance to question entities, ranked by their embedding similarity to the query, until the top-K chunks are selected. We call this approach Graph Retriever.
Additionally, as shown in step 4 of Figure 2, we prompt an LLM to identify the necessary propositions to answer the question from those retrieved in the Graph Retriever process, effectively using LLM reasoning capabilities to re-rank the selected propositions. Following this LLM-based re-ranking, we include the chunks corresponding to the LLM-identified propositions first, and then fall back to the Graph Retriever to select additional chunks until the top-K chunks are selected. We refer to this combined approach as Graph+LLM in Section 6.
@@ 6.1 KG Synthesis and Distillation Settings
We use Llama-3.1-70B-Instruct111https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct on 100K documents from BAAI/IndustryCorpus222https://huggingface.co/datasets/BAAI/IndustryCorpus for synthesizing KGs. We sample an equal number of documents from each of the following industry categories: politics, news, medicine, literature, finance, film & TV, computer science, automotive, technology, and education. We use the SentenceSplitter from the Llama-Index Liu (2022) framework to split documents into chunks, setting the chunk size to 256 tokens and chunk overlap to 0 tokens. We apply a filtering criterion based on the ROUGE-1 F1 score Lin (2004), setting a threshold of 0.70 to minimize the risk of hallucinations from decontextualization.
We perform the KG synthesis using VLLM Kwon et al. (2023) on 160 Intel® Gaudi 2 AI accelerators in the Intel® Tiber™ AI Cloud. We will publicly release our 100K generated document-KG pairs.
@@
We train the Meta-Llama-3-8B-Instruct model333https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) on 30K synthesized documents to directly generate corresponding KGs for entire input document. We use Optimum-Habana444https://github.com/huggingface/optimum-habana and the Transformers Wolf et al. (2020) Python library to train this distilled model on 8 Intel® Gaudi 2 AI accelerators in the Intel® Tiber™ AI Cloud. We employ a learning rate of 5e-5, a batch size of 32, and train for one epoch. We name our model .
| Dataset | KG Source | Triplets | Semantic Score | Triplet Coverage | F1 |
| MuSiQue | Llama-3-8b | 93855 | 0.8111 | 32.09 | 0.51 |
| Llama-3-70b | 102119 | 0.8346 | 40.34 | 0.56 | |
| 139376 | 0.8546 | 46.90 | 0.59 | ||
| 2wiki | Llama-3-8b | 41384 | 0.8281 | 43.39 | 0.56 |
| Llama-3-70b | 46100 | 0.8475 | 54.10 | 0.58 | |
| 68800 | 0.8693 | 58.27 | 0.59 | ||
| HotpotQA | Llama-3-8b | 76906 | 0.8343 | 41.79 | 0.58 |
| Llama-3-70b | 82948 | 0.8440 | 47.20 | 0.61 | |
| 123458 | 0.8693 | 55.26 | 0.64 |
| Dataset | KG Source | Retriever | Hits@2 | Hits@5 | Hits@10 | MRR | MAP |
| MuSiQue | None | Dense Retriever | 41.32 | 55.28 | 64.19 | 79.89 | 40.17 |
| None | Dense + LLM | 47.60 | 60.47 | 67.02 | 84.44 | 44.26 | |
| Llama-3-8b | Graph + LLM | 31.33 | 39.19 | 42.68 | 60.67 | 29.49 | |
| Llama-3-70b | Graph + LLM | 48.64 | 63.14 | 68.93 | 85.24 | 45.20 | |
| Graph + LLM | 53.35 | 67.93 | 72.78 | 87.41 | 48.04 | ||
| 2wiki | None | Dense Retrieval | 62.22 | 70.62 | 74.72 | 97.86 | 55.73 |
| None | Dense + LLM | 72.63 | 75.87 | 76.70 | 97.77 | 58.65 | |
| Llama-3-8b | Graph + LLM | 41.55 | 44.2 | 45.6 | 66.53 | 36.7 | |
| Llama-3-70b | Graph + LLM | 68.73 | 73.05 | 74.47 | 97.32 | 57.42 | |
| Graph + LLM | 73.15 | 77.5 | 78.57 | 98.74 | 59.91 | ||
| HotpotQA | None | Dense Retrieval | 66.55 | 81.80 | 89.45 | 91.98 | 60.68 |
| None | Dense + LLM | 83.10 | 90.05 | 92.10 | 96.79 | 67.58 | |
| Llama-3-8b | Graph + LLM | 50.65 | 55.5 | 57.45 | 73.72 | 45.06 | |
| Llama-3-70b | Graph + LLM | 79.10 | 90.35 | 93.75 | 93.27 | 65.78 | |
| Graph + LLM | 81.85 | 92.35 | 94.70 | 94.53 | 67.22 |
@@ 6.2 Evaluation Settings
We evaluate KGs extracted by our model on KG coverage, text chunk retrieval and QA tasks using 3 multi-hop reasoning datasets: MuSiQue, 2WikiMultiHopQA (2wiki) and HotpotQA. We follow the settings used in HippoRAG Gutiérrez et al. (2024), and use the exact same 1000 questions and candidate passages, including both supporting and distractor passages. For KG coverage evaluation, we generate proxy ground-truth triplets using the AzureOpenAI and GPT-4o models.
@@
Across all tasks, we compare KGs extracted by our against KGs extracted by two baseline models: Llama-3-8b and Llama-3-70b555We use the Instruct variants of both models throughout. For the retrieval and multihop QA tasks, we include the performance of the most widely used dense vector retrieval method, as well as a dense retriever combined with an LLM-based re-ranking approach. Lastly, for the multihop QA task, we also include results for a non-RAG system where the LLM utilizes its internal parametric knowledge to answer questions. We provide full experimental details in Appendix D.
@@
We evaluate our Distill-SynthKG model using three distinct indexing/ retrieval frameworks. In all frameworks, we employ LlamaIndex’s TreeSummarize response builder, powered by GPT-4o, to generate answers from the retrieved context. We only modify the query engine’s prompt by adding specific instructions to produce concise answers tailored to the requirements of the test datasets.
LlamaIndex: We use the KnowledgeGraphIndex module from LlamaIndex to build a hybrid KG index that combines keyword-based entity search with semantic similarity-based triplet search. This approach retrieves the top-K relevant text chunks along with the associated subgraph. Both the retrieved text chunks and the subgraph are then used to answer questions.
Chain-of-Triplet: We decompose a multi-hop question into chains of triplet sub-queries and retrieve the top 20 matching triplets from the knowledge graph for each sub-query. We then generate answers based on these triplets and the associated propositions. This approach allows us to directly assess the effectiveness of the extracted KGs in answering multi-hop questions. Refer to Appendix D.3.2 for more details of this method.
Graph+LLM: We use our proposition-entity graph retriever to retrieve the top 10 relevant chunks and 2-hop paths from the KG that include the question entity to generate the final answer.
@@ 7 Results
7.1 SynthKG: Triplets Density
In Figure 3, we compare the number of triplets generated per 100 words in documents of varying lengths using SynthKG versus a single-step LLM prompting approach. The analysis is conducted on a subset of 1,000 documents used for training Distill-SynthKG model. We observe that for the single-step prompting approach, the ratio of extracted relations decreases as document length increases, with the triplet density dropping by approximately 60% from 100-word documents to those of 1200 words. In contrast, for SynthKG, the density remains almost constant throughout, demonstrating its effectiveness in maintaining triplet generation across documents of varying lengths.
@@ 7.1 SynthKG: Triplets Density
| Graph + LLM | 0.320 | 0.459 | 0.440 | 0.544 | 0.539 | 0.706 | 0.433 | 0.569 | |
@@
| 0.695 | 0.792 | 0.820 | 0.936 | 0.584 | 0.433 | 0.569 | |
7.2 KG Coverage Results
Table 1 presents the results of the coverage evaluation of KGs generated by our model, along with two baseline models. We observe a consistent pattern across all three test datasets: the KGs generated by our model outperform both the untrained Llama-3-8b model of the same size and the much larger Llama-3-70b model across all metrics and datasets. Particularly noteworthy is that, compared to the KGs generated by Llama-3-70b, our model generates an additional 100,467 triplets across the combined three datasets, resulting in a 6.26% higher average absolute triplet coverage. These results underscore the superior quality of the KGs synthesized by SynthKG and demonstrate the effectiveness of distilling this multi-step process into a significantly smaller 8b model through fine-tuning.
7.3 Retrieval Results
We present the retrieval results in Table 2. The higher KG coverage of the model leads to consistently superior retrieval performance across all datasets and metrics, outperforming both the unfinetuned 8b-Instruct model of the same size and the much larger 70b-Instruct model. Specifically, the graph+LLM retriever using the KGs generated by our model achieves an average absolute improvement of 28.27 in hits@2 compared to the unfinetuned Llama-3-8b model and 3.96 in hits@2 over the larger Llama-3-70b model. Additionally, the graph+LLM retriever with our KG model delivers an average improvement of 12.75 in hits@2 over standard dense retrieval and 1.67 in hits@2 over the dense retriever with an LLM-based reranker.
7.4 Multi-hop QA Results
We present the results of the multi-hop QA task in Table 3, evaluating the general applicability of the KGs generated by our model across different retrieval and QA frameworks. A good model, along with its extracted KGs, should ideally demonstrate robustness across various retrieval and question-answering methodologies. The results indicate that under three distinct QA frameworks (LlamaIndex, Chain-of-Triplet, and Graph + LLM), our model’s generated KGs consistently outperform its unfinetuned 8B counterpart as well as the larger 70B model. In the Chain-of-Triplet framework, which relies exclusively on the KG for answering questions, our distilled model shows an absolute gain of +11.6% in EM accuracy compared to the Llama-3-8b model, while the improvement is +5.9% in the LlamaIndex framework. This further underscores the superior coverage of the KG extracted by our distilled model. In the Graph + LLM framework, our distilled model achieves the highest gain, with a +15.2% absolute increase in EM accuracy over the Llama-3-8b model, resulting in the best overall performance. Additionally, both our model and the Llama-3-70b model with the Graph + LLM retriever outperform the dense retriever as well as the dense retriever with LLM reranker, highlighting the importance of using our entity-proposition graph retriever in conjunction with the KG to achieve optimal results.
7.5 Ablations
Efficiency and Generalizability of Distill-SynthKG
In Table 4, we study three key important questions for developing Distill-SynthKG: 1. the efficiency of training Distill-SynthKG, 2. the effectiveness of various powerful LLMs for synthesizing training data, and 3. the generalizability of fine-tuning other smaller LLMs on synthesized data. To address these questions, we employ QLoRA Dettmers et al. (2023) fine-tuning on approximately 1,000 synthetic Document-KG pairs, generated using GPT-4o and Llama-3.1-70B-Instruct. We provide fine-tuning configurations in Section C.1. Additionally, to answer the third question, we fine-tune another well-known base LLM, Mistral-7B-Instruct-v0.3 Jiang et al. (2023).
@@ 7.5 Ablations
Benefits of KG in multi-hop QA
Along with text chunks, our KG provides both triplets (or relation paths between entities) and propositions, which can all be leveraged as context to answer questions. To determine the optimal combination of these inputs for multi-hop QA, we conduct an ablation study, as shown in Figure 4. Our findings indicate that adding either related propositions (+2 EM) or 2-hop paths (+2.7 EM) from the KG to the top-K chunks improves performance on multi-hop QA task. Notably, 2-hop paths perform slightly better, likely due to their ability to capture more complex relationships between entities by connecting multiple relations across two propositions. Lastly, combining both triplets and propositions did not result in further improvement, which may be due to redundant information being captured by both triplets and propositions.
KG-based Retrieval Methods
Our KG supports multiple retrieval methods: (1) retrieval based on the similarity between triplets and questions, (2) retrieval based on the similarity between propositions and questions, and (3) graph-based retrieval, which ranks chunks by integrating entities and propositions. Optionally, all methods can incorporate an LLM for reranking. In Figure 5, we compare three retrieval methods utilizing KGs. Our findings indicate that propositions outperform triplets, achieving an increase of +0.89 in Hits@10, which is expected, as propositions offer more detailed information than triplets. The graph retrieval method demonstrates even greater performance by integrating both entities and propositions, yielding a +2.50 higher Hits@10 score compared to propositions alone. Secondly, for all three retrieval types, incorporating an LLM to re-rank propositions and triplets significantly enhances retrieval performance, with an increase of +3.59 in Hits@10 for the best performing graph method.
8 Conclusion
In this work, we introduced SynthKG, a novel approach for synthesizing high-coverage KG training data using LLMs. Leveraging this synthesized data, we proposed Distill-SynthKG, an efficient model that distills the multi-step KG construction process into a single inference step. Our experiments demonstrated that Distill-SynthKG significantly improves KG coverage, retrieval accuracy, and QA performance across multiple datasets, outperforming a model nearly eight times its size. These results validate the effectiveness of Distill-SynthKG, highlighting its potential for scalable, ontology-free KG construction and its application in RAG tasks.
9 Limitations
The documents used for synthesizing KGs in this study were limited to the English language, which has been widely studied in existing NLP research. Additional work is needed to investigate the construction of KGs from documents in other languages. While we investigated synthetic data generation using two strong foundational LLMs (Llama-3-70b and GPT-4o), the use of other LLMs with SynthKG may yield different results than those reported in our study. Similarly, our distillation experiments were limited to two popular LLMs (Llama-3-8b and Mistral-7b) which we believe are representative of the capabilities of LLMs of similar size. Finally, our proposed benchmarks for evaluating KG coverage rely on automated question decomposition and triplet extraction using GPT-4o, which introduces the possibility of errors or omissions (see Appendix B.2 for human evaluation results).
Appendix A LLM Prompts for SynthKG
We use prompts Figure 6, Figure 7 and Figure 8 for decontextualization, entity extraction and relation extraction respectively within SynthKG. We also provide an example of decontextualized chunk in Figure 9.
Previous paragraph from Document:

### v2

@@ 5 Proposition-Entity Graph Retriever
We introduce a retrieval framework that leverages the structure of our synthesized KGs, in particular the proposition–entity bipartite graph (Figure˜2). Unlike prior graph-based retrieval methods that operate over sparse triplets, our retriever treats propositions as the primary retrieval units. Because propositions are rich, self-contained text statements, they serve as higher-quality candidates for RAG while the bipartite topology provides explicit links through shared entities, promoting logical connectivity across retrieved evidence. This design is made possible by the consistent KG structure produced by our data-centric pipeline.
Given a question, we first retrieve the top-M most relevant propositions from the KG using embedding similarity, narrowing the search space to a smaller subset of relevant information. This initial semantic retrieval is crucial because propositions, unlike raw triplets, contain sufficient context for meaningful similarity computation. In step 2, we construct a sub-graph consisting of these propositions and their linked entities, capturing the relations among the retrieved propositions. In step 3, we traverse the sub-graph starting from the entities mentioned in the question, selecting only propositions within their N-hop neighborhood. This graph-constrained filtering addresses a key challenge in multi-hop reasoning: distinguishing between semantically related but logically disconnected information. For instance, propositions about “Washington” the president versus “Washington” the state may have high embedding similarity but different graph neighborhoods. We then include text chunks corresponding to the selected propositions within N-hop distance to question entities, ranked by their embedding similarity to the query, until the top-K chunks are selected. We call this approach Graph Retriever.
Additionally, as shown in Step 4 of Figure˜2, we augment retrieval with LLM-based re-ranking. Given the candidate propositions returned by the Graph Retriever, we prompt an LLM to select the propositions required to answer the question, using its reasoning ability to re-rank the evidence. This is feasible because propositions are interpretable text units that an LLM can directly judge for relevance, unlike isolated triplets.
After re-ranking, we prioritize the chunks associated with the LLM-selected propositions, and then fall back to the Graph Retriever to fill the remaining budget until the top-K chunks are selected. We refer to this hybrid method as Graph+LLM in Section˜6.
The design of this retrieval framework is tightly coupled with our data synthesis approach: the consistent entity naming from decontextualization ensures reliable graph traversal, while the proposition generation provides semantically rich retrieval units.
@@ 6 Experiments
Our experiments are designed to validate the data-centric approach across three dimensions: (1) the quality of synthetic training data, (2) the effectiveness of distillation, and (3) the performance of downstream applications. We address the following research questions: (RQ1) Does the multi-step SynthKG pipeline produce higher-quality KGs than direct prompting? (RQ2) Can an 8B model achieve large-model performance through distillation on synthetic data? (RQ3) How does data scale affect the performance of distilled models? (RQ4) Does improved KG quality translate to better retrieval and QA performance?
@@ 6.1 KG Synthesis and Distillation Settings
To create a large-scale training resource, we use Llama-3.1-70b-Instruct (AI@Meta, 2024) to synthesize KGs from 100K documents from IndustryCorpus (by BAAI). We ensure domain diversity by sampling equally from ten categories: politics, news, medicine, literature, finance, film & TV, computer science, automotive, technology, and education. We use the SentenceSplitter from the Llama-Index (Liu, 2022) framework to split documents into chunks, setting the chunk size to 256 tokens and chunk overlap to 0 tokens. We apply a filtering criterion based on the ROUGE-1 F1 score (Lin, 2004), setting a threshold of 0.70 to minimize the risk of hallucinations from decontextualization. We perform the KG synthesis using VLLM (Kwon et al., 2023) on 160 Intel® Gaudi 2 AI accelerators in the Intel® Tiber™ AI Cloud. Our 100K generated document-KG pairs represent the first large-scale training resource for ontology-free KG construction and will be publicly released.
@@
We validate that synthetic data enables effective distillation by training Meta-Llama-3-8b-Instruct (AI@Meta, 2024) on 30K synthesized documents. The model learns to directly generate corresponding KGs for entire input documents in a single pass. Training uses 8 Intel® Gaudi 2 AI accelerators with a learning rate of 5e-5, batch size of 32, for one epoch. We name our model Distill-SynthKG and refer to it subsequently as D-SynthKG-8b.
@@ 6.2 Evaluation Settings
We evaluate on three multi-hop reasoning benchmarks that require integrating information across multiple facts: MuSiQue, 2WikiMultiHopQA (2Wiki), and HotpotQA. These datasets are well suited for KG evaluation because answering their questions depends on structured, multi-fact reasoning, which is precisely the behavior KGs are intended to support. Following the experimental protocol of HippoRAG (gutiérrez2024hipporag), we use the same 1,000 questions and the same candidate passage sets, which include both supporting and distractor passages, to enable a fair comparison. For KG coverage evaluation, we generate proxy ground-truth triplets with GPT-4o as described in Section 4.
@@
We compare against multiple baseline categories for comprehensive evaluations:
Direct prompting baselines: Llama-3-8b and Llama-3-70b111We use the Instruct variants of both models throughout for single-step KG extraction
Multi-step baselines: Full SynthKG pipeline with Llama-3-8b (SynthKG-8b) and Llama-3-70b (SynthKG-70b) to assess distillation effectiveness
Retrieval baselines: Standard dense vector retrieval and a two-stage approach combining dense retrieval with LLM-based re-ranking (Dense+LLM) to establish non-KG performance benchmarks
KG-based retrieval: Retrieval using GPT-4o-based KGs
KG-RAG systems: GraphRAG and HippoRAG using GPT-4o-extracted KGs, representing state-of-the-art KG-based approaches222To improve GraphRAG performance, we append the instruction: “Only provide the answer without any context. For yes/no questions, just mention yes or no. Do not cite data sources.” at the end of each query. For GraphRAG, we report results using the local and drift modes, which yield the best performance; the global mode is excluded. For HippoRAG, we use GPT-4o for KG construction and apply our query synthesizer to the retrieved text chunks to generate the final answer, ensuring a fair comparison.
Closed-book QA: LLM using only parametric knowledge, establishing the lower bound
We provide full experimental details including hyperparameters in Appendix˜F.
@@
To demonstrate the versatility of our KGs, we evaluate using three distinct retrieval frameworks using LlamaIndex’s TreeSummarize with GPT-4o for answer generation:
LlamaIndex: Standard KG-based retrieval using LlamaIndex’s KnowledgeGraphIndex with hybrid keyword and semantic search
Chain-of-Triplet: Direct KG reasoning by decomposing questions into sub-queries and retrieving matching triplets (details in Appendix F.3.2)
Graph+LLM: Our proposition-entity graph retriever with LLM re-ranking, leveraging the unique structure of our synthesized KGs
@@ 7 Results
| KG Source | MuSiQue | 2wiki | HotpotQA | |||||||||
| Triplets | Semantic | Coverage | F1 | Triplets | Semantic | Coverage | F1 | Triplets | Semantic | Coverage | F1 | |
| Llama-3-8b | 93855 | 0.8111 | 32.09 | 0.51 | 41384 | 0.8281 | 43.39 | 0.56 | 76906 | 0.8343 | 41.79 | 0.58 |
| SynthKG-8b | 125197 | 0.8341 | 38.84 | 0.55 | 56178 | 0.8275 | 44.56 | 0.54 | 108031 | 0.8448 | 47.72 | 0.60 |
| Llama-3-70b | 102119 | 0.8346 | 40.34 | 0.56 | 46100 | 0.8475 | 54.10 | 0.58 | 82948 | 0.8440 | 47.20 | 0.61 |
| SynthKG-70b | 140527 | 0.8559 | 47.18 | 0.59 | 71305 | 0.8778 | 63.30 | 0.61 | 124460 | 0.8633 | 54.54 | 0.63 |
| D-SynthKG-8b | 139376 | 0.8546 | 46.90 | 0.59 | 68800 | 0.8693 | 58.27 | 0.59 | 123458 | 0.8693 | 55.26 | 0.64 |
We present experimental results that validate our data-centric approach to KG construction, directly addressing the research questions posed in Section 6.
7.1 KG Coverage Results (RQ1: Multi-step vs Direct Prompting)
The multi-step SynthKG pipeline consistently generates more triplets and achieves higher coverage than the commonly used single-step LLM prompting approach across all three datasets, for both LLaMA-3-8b and 70b models (Table˜1). This validates our hypothesis that decomposing KG construction into systematic steps prevents the information loss inherent in single-pass processing. Furthermore, our D-SynthKG-8b model outperforms the untrained Llama-3-8b, Llama-3-70b, and SynthKG-8b baselines, demonstrating the benefit of distilling the SynthKG pipeline using Llama-3-70b as the teacher. Remarkably, D-SynthKG-8b is also highly competitive with SynthKG-70b, despite being approximately 8 smaller and relying on a single-step inference process. These results underscore that high-quality training data can bridge the gap between model sizes. As our KG coverage metric emphasizes recall of critical information, we also verified precision through manual inspection. Among 150 randomly sampled triplets from D-SynthKG-8b’s predictions, only 4 were incorrect and 5 meaningless, indicating that the generated triplets largely align with the source content while maintaining high coverage.
7.2 Retrieval (RQ2 & RQ4: Distillation Effectiveness and Downstream Impact)
| KG Source | Retriever | MuSiQue | 2wiki | HotpotQA | |||||||||
| Hits@2 | Hits@10 | MRR | MAP | Hits@2 | Hits@10 | MRR | MAP | Hits@2 | Hits@10 | MRR | MAP | ||
| None | Dense | 41.32 | 64.19 | 79.89 | 40.17 | 62.22 | 74.72 | 97.86 | 55.73 | 66.55 | 89.45 | 91.98 | 60.68 |
| None | Dense+LLM | 47.60 | 67.02 | 84.44 | 44.26 | 72.63 | 76.70 | 97.77 | 58.65 | 83.10 | 92.10 | 96.79 | 67.58 |
| Llama-3-8b | Graph + LLM | 31.33 | 42.68 | 60.67 | 29.49 | 41.55 | 45.60 | 66.53 | 36.70 | 50.65 | 57.45 | 73.72 | 45.06 |
| SynthKG-8b | Graph + LLM | 50.62 | 65.17 | 86.65 | 45.43 | 65.25 | 69.65 | 95.54 | 54.79 | 76.55 | 86.35 | 92.69 | 63.44 |
| Llama-3-70b | Graph + LLM | 48.64 | 68.93 | 85.24 | 45.20 | 68.73 | 74.47 | 97.32 | 57.42 | 79.10 | 93.75 | 93.27 | 65.78 |
| SynthKG-70b | Graph + LLM | 53.70 | 72.23 | 88.81 | 48.32 | 73.23 | 78.80 | 98.80 | 60.09 | 81.90 | 94.40 | 94.62 | 66.93 |
| D-SynthKG-8b | Graph + LLM | 53.35 | 72.78 | 87.41 | 48.04 | 73.15 | 78.57 | 98.74 | 59.91 | 81.85 | 94.70 | 94.53 | 67.22 |
| GPT-4o | Graph + LLM | 53.90 | 70.38 | 90.46 | 48.66 | 74.35 | 79.25 | 99.02 | 60.52 | 82.90 | 94.95 | 93.98 | 67.15 |
Our D-SynthKG-8b model improves Hits@2 by an average of 28.27 points over the pre-trained Llama-3-8b, 5.31 points over SynthKG-8b, and 3.96 points over the larger Llama-3-70b (Table˜2). These gains suggest that training on synthetic data produces a qualitative shift in capability rather than a marginal tuning effect. Notably, D-SynthKG-8b remains highly competitive with the full SynthKG-70b pipeline, despite being much smaller and requiring only single-step inference, indicating that the multi-step procedure can be effectively internalized through distillation. It also performs comparably to GPT-4o (see Appendix B.2 for details).
In addition, our Graph+LLM retriever improves Hits@2 by 12.75 points over standard dense retrieval and by 1.67 points over dense retrieval augmented with an LLM re-ranker, demonstrating that structured KGs enable more effective retrieval strategies.
@@ 7.2 Retrieval (RQ2 & RQ4: Distillation Effectiveness and Downstream Impact)
| D-SynthKG-8b | Graph + LLM | 0.320 | 0.459 | 0.440 | 0.544 | 0.539 | 0.706 | 0.433 | 0.569 |
7.3 Multi-hop QA Results (RQ4: End-to-End Performance)
D-SynthKG-8b delivers the best overall results across all three frameworks: LlamaIndex, Chain-of-Triplet, and Graph+LLM, surpassing both Llama-3-8b and the larger Llama-3-70b (Table˜3). The consistency of these gains across diverse RAG pipelines highlights the robustness and broad applicability of our synthesized KGs. The largest improvement appears in the Graph+LLM setting, where D-SynthKG-8b achieves a +15.2% absolute EM gain over Llama-3-8b and a +2.0% gain over Llama-3-70b, yielding the best overall performance. Notably, it also outperforms GraphRAG and HippoRAG, two strong KG-based RAG systems built on KGs generated by GPT-4o, underscoring that our data-centric approach enables a smaller model to compete with substantially larger systems in practical applications.
7.4 Analysis (RQ3: Scaling and Design Validation)
How effective is multi-step SynthKG in processing documents of increasing length?
We compare our multi-step SynthKG framework with a single-step LLM prompting approach by examining the number of triples generated per 100 words for documents of varying lengths (Figure˜4, Appendix B). For the single-step method, the proportion of extracted relations decreases as document length increases, with triple density dropping by about 60% when moving from 100-word documents to 1,200-word documents. In contrast, SynthKG’s triple density remains nearly constant across all lengths. This validates our chunking strategy and demonstrates why the multi-step approach is essential for creating consistent training data.
What is the optimal retrieval source?
Our KGs support multiple retrieval strategies. In Figure˜3(a) in Appendix B, we compare retrieval using triples, propositions, and the full graph structure. Proposition retrieval outperforms triples (+0.89 Hits@10) due to richer context, while graph-based retrieval achieves the best performance (+2.50 over propositions). This validates our design choice of including propositions as first-class components of the KG, not just intermediate representations.
How can our KG improve RAG beyond retrieval?
Our ablation study (Figure˜3(b) in Appendix B) shows that enriching retrieved context with structured signals improves QA performance. Adding propositions (+2 EM) or 2-hop paths (+2.7 EM) to retrieved chunks improves accuracy, with 2-hop paths offering slightly better gains due to their ability to capture complex relationships. This demonstrates that our KGs provide value beyond simple retrieval, enabling structured reasoning.
8 Conclusion
We presented SynthKG, a data synthesis pipeline that generates high-quality document-KG pairs, and Distill-SynthKG, which distills this multi-step process into an efficient single-step model. Our experiments demonstrate that an 8B model trained on synthetic data matches or exceeds the performance of models eight times larger across KG quality, retrieval, and QA tasks. This work shifts the paradigm from scaling models to generating training data, showing that structured extraction capabilities can be transferred through synthetic examples rather than parameter count. We release our 100K document-KG pairs to enable future research in data-centric approaches to knowledge extraction.
Reproducibility Statement
To ensure reproducibility, we provide detailed experimental settings and hyperparameters in Section 6 and Appendix F. All prompts used in the SynthKG pipeline are included in Appendix C. The 100K synthetic document-KG pairs dataset will be made publicly available. The core implementation code, including the KG synthesis pipeline and evaluation metrics, will be released to facilitate the reproduction of our results.
Appendix A Appendix
Appendix B Additional Analyses
B.1 Efficiency and Generalizability of Distill-SynthKG
@@
| D-SynthKG-8b | 0.695 | 0.792 | 0.820 | 0.936 | 0.584 | 0.433 | 0.569 |
In Table 4, we study three key important questions for developing Distill-SynthKG: 1. the efficiency of training Distill-SynthKG, 2. the effectiveness of various powerful LLMs for synthesizing training data, and 3. the generalizability of fine-tuning other smaller LLMs on synthesized data. To address these questions, we employ QLoRA (Dettmers et al., 2023) fine-tuning on approximately 1,000 synthetic Document-KG pairs, generated using GPT-4o and Llama-3.1-70b-Instruct. We provide fine-tuning configurations in Section˜E.1. Additionally, to answer the third question, we fine-tune another well-known base LLM, Mistral-7b-Instruct-v0.3 (Jiang et al., 2023).
@@
| Dataset | Model | Hits@2 | Hits@10 | MAP | MRR |
| 2Wiki | GPT-4o (OpenAI) | 74.35 | 79.25 | 60.52 | 99.02 |
| D-SynthKG-8b | 73.15 | 78.57 | 59.91 | 98.74 | |
| HotpotQA | GPT-4o (OpenAI) | 82.90 | 94.95 | 67.15 | 93.98 |
| D-SynthKG-8b | 81.85 | 94.70 | 67.22 | 94.53 | |
| MuSiQue | GPT-4o (OpenAI) | 53.90 | 70.38 | 48.66 | 90.46 |
| D-SynthKG-8b | 53.35 | 72.78 | 48.04 | 87.41 |
B.2 Comparison of Retrieval Performance with Proprietary Foundation Models
We include additional comparisons between our distilled model D-SynthKG-8b and a state-of-the-art proprietary foundation model, GPT-4o from OpenAI. We evaluate retrieval performance on three multihop QA benchmarks: 2Wiki, HotpotQA, and MuSiQue. We present the results in Table˜5. The results show that D-SynthKG-8b achieves retrieval performance that is highly competitive with GPT-4o across all datasets. Notably, D-SynthKG-8b yields a slightly higher average Hits@10 (82.02 vs. 81.52), despite being significantly more cost-efficient. GPT-4o incurs $2.50 per million input tokens and $10 per million output tokens, while D-SynthKG-8b operates at only $0.20 per million tokens (input or output), representing roughly 3% of the inference cost. These findings highlight the practical advantages of our approach in cost-sensitive deployment scenarios.
B.3 Entity Distribution Analysis
When analyzing entity references across document chunks, we find that the overall average chunk distance per entity is 0.9. This metric represents the average number of chunks between an entity mention and its most recent previous mention. Further analysis reveals that 80.03% of entities have their most recent mention within a single paragraph, indicating that the majority of entity references are relatively localized within the document.
These findings support our design decision to only reference the immediately preceding chunk during decontextualization, as this approach effectively balances computational efficiency with adequate contextual information.
Appendix C LLM Prompts for SynthKG
We use prompts Figure˜5, Figure˜6 and Figure˜7 for decontextualization, entity extraction and relation extraction respectively within SynthKG. We also provide an example of decontextualized chunk in Figure˜8.